Which coordinate notation represent the translation of P(-2, 5) to P’ (9, -12)?
1
Expert's answer
2019-09-06T09:54:16-0400

Solution. A Translation Vector is a vector that gives the length and direction of a particular translation. Vectors in the Cartesian plane can be written (x,y) which means a translation of x units horizontally and y units vertically. Find the coordinates of the translation vector.


"(9-(-2),-12-5)=(11,-17)"

On the other hand, we can write the coordinate transformation equations


"x'=x+11, y'=y-17"

Answer.

"x'=x+11, y'=y-17"